An instructor has given a short quiz consisting of two parts. For a randomly selected student, let X = the number of points earned on the first part and Y = the number of points earned on the second part. Suppose that the joint pmf of X and Y is given in the accompanying table.

y

p(x, y)

     
  0      5      10      15   
x 0 0.01 0.06 0.02 0.10
5 0.04 0.14 0.20 0.10
10 0.01 0.15 0.16 0.01

(a) If the score recorded in the grade book is the total number of points earned on the two parts, what is the expected recorded score E(X + Y)? (Enter your answer to one decimal place.)


(b) If the maximum of the two scores is recorded, what is the expected recorded score? (Enter your answer to two decimal places.)
Answer #1

a) The value of X + Y is computed for each combination of X, Y here. The expected value for X + Y then is computed here as:

E(X + Y) = \sum (x + y)P(X = x, Y = y)

p(x,y)   0      5      10      15   
0 0.01 0.06 0.02 0.1
5 0.04 0.14 0.2 0.1
10 0.01 0.15 0.16 0.01
x + y   0      5      10      15   
0 0 5 10 15
5 5 10 15 20
10 10 15 20 25
E(X + Y) 0 5 10 15
0 0 0.3 0.2 1.5
5 0.2 1.4 3 2
10 0.1 2.25 3.2 0.25

We add the values in the last table here to get the expected value as:

E(X + Y) = \sum (x + y)P(X = x, Y = y) = 14.4

Therefore 14.4 is the required expected value here.

b) For the maximum of the two scores, we calculate the expected score here as:

Max(X, Y) 0 5 10 15
0 0 5 10 15
5 5 5 10 15
10 10 10 10 15
E(max(X, Y)) 0 5 10 15
0 0 0.3 0.2 1.5
5 0.2 0.7 2 1.5
10 0.1 1.5 1.6 0.15

E(max(X, Y)) = \sum Max(X, Y )P(X = x, Y = y) = 9.75

Therefore 9.75 is the required expected value here.